# Functional normalizing flow for statistical inverse problems of partial differential equations

### v1

@@ 4.2.1 Functional Planar Flow
Figure 8: The comparison of the approximate posterior obtained by functional planar flow with 24 layers and the background truth. (a)(d): The background truth of $u$. (b): The mean of approximate posterior obtained by functional planar flow with 24 layers. (c):The variance function of approximate posterior obtained by functional planar flow with 24 layers. (e): The mean function of posterior obtained by pCN algorithm. (f): The variance function of posterior obtained by pCN algorithm.
To further support the effectiveness of the NF-iVI method in estimating the mean function, we provide numerical evidence. The relative error between the mean of the approximate posterior $u_{N}^{*}$ and the mean of the posterior $u_{p}^{*}$ is given by:
@@
The estimated value is small, indicating the effectiveness of our algorithm.
Combining the visual evidence from Figure 8 and the numerical evidence provided by the relative error in (4.8), we conclude that the NF-iVI method effectively approximates the mean of the posterior.
Secondly, to further explore the approximate posterior, we analyze the covariance functions. In subfigure (c) of Figure 8, we present the point-wise variance field of the posterior measure obtained from the NF-iVI method. For comparison, subfigure (f) of Figure 8 depicts the point-wise variance field estimated using the pCN method. The similarity between these two plots indicates the effectiveness of the NF-iVI method in quantifying the uncertainties of the parameter $u$.
We also compare the covariance functions generated by these two methods. For simplicity, we compute these quantities at the mesh points and present the results in Figure 9. In all subfigures of Figure 9, the estimated covariance of posteriors obtained by the pCN and NF-iVI methods are represented by a blue solid line and a red dashed line, respectively. In subfigure (a) of Figure 9, we show the variance function calculated at all mesh points, i.e., $\{\mathrm{var}_{u}(x_{i})\}_{i=1}^{N_{g}}~{}(N_{g}$ is the number of mesh points, which is 441 in this problem). In subfigures (c) and (e) of Figure 9, we show the covariance functions calculated on the pairs of points $\left\{(x_{i},x_{i+40})\right\}_{i=1}^{N_{g}-40}$ and $\{(x_{i},x_{i+80})\}_{i=1}^{N_{g}-80}$, respectively.
To numerically compare the variance and covariance functions generated by the two methods, we present the relative errors in Table 7. In the table, the notation $\bm{c}$ means the total relative error, and the notation $\{c(x_{i},x_{i+k})\}^{N_{g}-k}_{i=1}$ means the covariance function values on the pair of mesh points $\{(x_{i},x_{i+k})\}^{N_{g}-k}_{i=1}$ ($k=0,40,80$). The numbers below this notation are the relative errors between the vectors obtained by pCN method and NF-iVI, respectively. The small relative errors indicate that the posterior covariance functions obtained by the NF-iVI and pCN sampling methods are quantitatively similar. Combining the visual evidence from Figures 8 and 9, and the numerical evidence from Table 7, we conclude that the NF-iVI method effectively approximates the covariance function of the posterior.
planar $0.12252$ $0.03424$ $0.11531$ $0.16419$
(a) Variance
(b) Covariance
(c) Covariance
Figure 9: The estimated variance and covariance functions obtained by the pCN algorithm (blue solid line), and NF-iVI algorithm (red dashed line). (a): The covariance function $c(x,y)$ on all the mesh point pairs $\{(x_{i},x_{i})\}^{N_{g}}_{i=1}$; (b): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+40})\}^{N_{g}-40}_{i=1}$; (c): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+80})\}^{N_{g}-80}_{i=1}$.
To demonstrate the discrete invariance of functional planar flow, we assess the impact of discretization by employing different grid sizes: $15\times 15$, $20\times 20$, $30\times 30$, $40\times 40$, and $50\times 50$ points in a regular mesh. We calculate the $L^{2}$-errors between the mean of the approximate posterior and the true value under these varying discrete levels. As presented in Table 8, the $L^{2}$-errors remains consistent across different discrete levels. Additionally, Table 9 compares the values of the covariance function of the approximate posterior for various discrete levels, revealing a similar pattern. These results collectively support the claim that the approximate posteriors obtained by our algorithm are invariant to different discretizations.
Discreate Level $15\times 15$ $20\times 20$ $30\times 30$ $40\times 40$ $50\times 50$
$L^{2}$-Error $0.009620$ $0.008446$ $0.007645$ $0.007913$ $0.008027$
Table 9: The values of covariance function, where $x_{1}=(0.3,0.3)$, $x_{2}=(0.5,0.5)$, $x_{3}=(0.7,0.7)$.
Discreate Level $15\times 15$ $20\times 20$ $30\times 30$ $40\times 40$ $50\times 50$
$c(x_{1},x_{1})$ $0.015079$ $0.016000$ $0.014736$ $0.015205$ $0.015848$
$c(x_{1},x_{2})$ $-0.004403$ $-0.004122$ $-0.004202$ $-0.004557$ $-0.455003$
$c(x_{1},x_{3})$ $0.004367$ $0.004269$ $0.004529$ $0.004891$ $0.004290$
$c(x_{2},x_{2})$ $0.015369$ $0.014359$ $0.013761$ $0.014083$ $0.013925$
4.2.2 Functional Sylvester Flow
As discussed in Subsection 2.4, the functional Sylvester flow is anticipated to exhibit superior approximation capabilities compared to the functional planar flow. To demonstrate this, we consider the same inverse problem defined in (4.7). In this experiment, we set the dimension of $\text{Im}(\mathcal{F}_{n})$ in the functional Sylvester flow to 20 ($M=20$ in Subsection 2.4). By employing only five layers of functional Sylvester flow, we achieve satisfactory results in approximating the posterior. In contrast, using five layers of functional planar flow would be insufficient to achieve similar performance.
We utilize five layers for both functional Sylvester flow and functional planar flow. Following Algorithm 1, all of the functional normalizing flow models are trained for $5000$ iterations $(K=5000)$, using 30 samples $(N=30)$ in each iteration, with an initial learning rate of $\alpha_{0}=0.01$. The learning rate follows a step decay schedule with a multiplicative factor of $\tau=0.8$ after every 500 iterations.
In subfigures (a) and (d) of Figure 10, we present the ground truth. Subfigures (b), (c), (e), and (f) of Figure 10 depict the mean function and the estimated point-wise variance field of the approximate posterior obtained from the 5-layer functional Sylvester flow. We observe that these results are consistent with those of the pCN algorithm shown in subfigures (e) and (f) of Figure 8, demonstrating the effectiveness of our approach.
@@ 4.2.2 Functional Sylvester Flow
(b) Estimated mean
(c) Estimated variance
@@
(e) Estimated mean
(f) Estimated variance
Figure 10: The comparison of approximation ability between functional planar flow and functional Sylvester flow. We use 5 layers for both flow models to approximate the posterior. (a)(d): The background truth of $u$. (b): The mean of approximate posterior obtained by functional Sylvester flow. (c): The variance function of approximate posterior obtained by functional Sylvester flow. (e): The mean function of approximate posterior obtained by functional planar flow. (c):The variance function of approximate posterior obtained by functional planar flow.
We also compare the variance and covariance functions generated by these two methods. We compute these quantities at the mesh points and present the results in Figure 11. In all subfigures of Figure 11, the estimates obtained by the pCN, functional planar flow, and functional Sylvester flow are drawn in blue solid line, orange solid line, and red dashed line, respectively. In subfigures (a) and (d) of Figure 11, we show the variance function calculated at all mesh points, i.e., $\{\mathrm{var}{u}(x_{i})\}_{i=1}^{N_{g}}~{}(N_{g}=441$ being the number of mesh points). In subfigures (b) and (e) of Figure 11, we show the covariance functions calculated on pairs of points $\left\{(x_{i},x_{i+40})\right\}_{i=1}^{N_{g}-40}$. Finally, in subfigures (c) and (f) of Figure 11, we show the covariance functions calculated on pairs of points $\{(x_{i},x_{i+80})\}_{i=1}^{N_{g}-80}$.
To numerically compare the variance and covariance functions, we employ the 5-layer functional planar flow and the 5-layer functional Sylvester flow to approximate the posterior. The relative errors between the variance functions and the covariance functions are presented in Table 10. As depicted in the table, the relative errors of functional Sylvester flow are consistently smaller than those of functional planar flow. Moreover, the variance and covariance functions obtained by the pCN method and functional Sylvester flow demonstrate quantitative similarity. This indicates that the 5-layer functional Sylvester flow exhibits superior performance in approximating the covariance function of the posterior compared to the 5-layer functional planar flow.
planar $206.804$ $155.946$ $161.661$ $174.727$
Sylvester $0.11846$ $0.05066$ $0.08278$ $0.15118$
The comparisons presented above highlight the limitations of functional planar flow in achieving accurate approximations with a small number of layers. In contrast, functional Sylvester flow exhibits superior approximation capabilities.
(a) Variance
(b) Covariance
(c) Covariance
diag
(d) Variance
(e) Covariance
(f) Covariance
Figure 11: The estimated variance and covariance functions obtained by the pCN method (blue solid line), planar NF-iVI method (green dashed line) and Sylvester NF-iVI (red dashed line). (a)(d): The covariance function $c(x,y)$ on all the mesh point pairs $\{(x_{i},x_{i})\}^{N_{g}}_{i=1}$; (b)(e): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+40})\}^{N_{g}-40}_{i=1}$; (c)(f): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+80})\}^{N_{g}-80}_{i=1}$.
Next, we demonstrate the discrete invariance of functional Sylvester flow. We calculate the $L^{2}$-errors between the mean of the approximate posterior and the true value under different discrete levels. As presented in Table 11, the $L^{2}$-errors remains consistent across different discrete levels. Additionally, Table 12 compares the values of the covariance function of the approximate posterior for various discrete levels, revealing a similar pattern. These results collectively support the claim that the approximate posteriors obtained by functional Sylvester flow are invariant to different discretizations.
Discreate Level $15\times 15$ $20\times 20$ $30\times 30$ $40\times 40$ $50\times 50$
$L^{2}$-Error $0.005291$ $0.005371$ $0.005293$ $0.005355$ $0.005166$
Table 12: The values of covariance function, where $x_{1}=(0.3,0.3)$, $x_{2}=(0.5,0.5)$, $x_{3}=(0.7,0.7)$.
Discreate Level $15\times 15$ $20\times 20$ $30\times 30$ $40\times 40$ $50\times 50$
$c(x_{1},x_{1})$ $0.015359$ $0.014409$ $0.014562$ $0.015302$ $0.014819$
$c(x_{1},x_{2})$ $-0.004174$ $-0.004598$ $-0.004539$ $-0.004566$ $-0.004515$
$c(x_{1},x_{3})$ $0.002194$ $0.001694$ $0.002079$ $0.001887$ $0.001986$
$c(x_{2},x_{2})$ $0.013898$ $0.013539$ $0.013355$ $0.013778$ $0.013336$
@@ 5 Numerical Examples of Conditional Functional Normalizing Flow
In this section, we present the results of the conditional functional normalizing flow, as detailed in Subsection 3. For a given measurement information pair $(\bm{x}^{*},\bm{d}^{*})$, the trained network $\mathcal{N}_{\lambda}(\bm{v})$ can efficiently provides a rough estimate of its corresponding posterior distribution $\mu_{(\bm{x}^{*},\bm{d}^{*})}$. Additionally, for a specific measurement information pair, we can refine the results through a retraining process.
@@ 5.1 Conditional Functional Normalizing Flow
In this subsection, we demonstrate the process of training the conditional network and obtaining an initial rough approximation of the posterior distribution directly from the trained network. We focus on the inverse problem of estimating the permeability distribution in a porous medium from a discrete set of pressure measurements, as studied in [11]. Considering the steady-state Darcy flow equation (4.6), the primary objective of the inverse problem is to infer the possible values of the parameter $u$ based on the measurement data $\bm{d}=(w(x^{1}),w(x^{2}),\ldots,w(x^{\bm{N}_{d}}))$ at the measurement points $\bm{x}=(x^{1},x^{2},\ldots,x^{\bm{N}_{d}})$ of the solution $w$. To approximate the posterior distribution, we employ a 5-layer functional Sylvester flow similar to the approach described in Subsection 4.2. To effectively train the conditional network, we generate the training dataset according to the following process:
@@
    Draw a large number of samples from the prior $\mu_{0}$, denoted as $\{u_{1},u_{2},\ldots,u_{N_{train}}\}$.
@@
(a) Truth
(b) CNF-iVI
(c) pCN
@@
Figure 12: The comparison of the approximate posterior obtained by CNF-iVI method and the pCN method. The estimated variance and covariance functions obtained by the pCN algorithm are drawn in blue solid line, and by CNF-iVI algorithm the red dashed line. (a): The background truth of $u$. (b): The mean of approximate posterior obtained by CNF-iVI method. (c): The mean of posterior obtained by pCN method. (d): The covariance function $c(x,y)$ on all the mesh point pairs $\{(x_{i},x_{i})\}^{N_{g}}_{i=1}$; (e): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+40})\}^{N_{g}-40}_{i=1}$; (f): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+80})\}^{N_{g}-80}_{i=1}$.
Here, we briefly describe the symbols used in this context. The vector $\bm{x}=(x^{1},x^{2},\ldots,x^{\bm{N}_{d}})$ represents the measurement points, where each $x^{i}\in\Omega$ corresponds to a distinct measurement point. The notation $\bm{x}_{j}$ refers to different measurement points associated with different measurement pairs $(\bm{x}_{j},\bm{d}_{j})$. For convenience, we select the measurement points $\bm{x}_{1},\bm{x}_{2},\ldots,\bm{x}_{N_{\text{train}}}$ within a bounded domain $\Omega$, which is defined as the region $(0,1)^{2}$. These points are uniformly distributed across this domain, with coordinates given by $\{(\frac{i}{21},\frac{j}{21})\}_{i,j=1}^{20}$.
Next, we introduce the training dataset $D_{train}$ into Algorithm 2 for training. We set $M=10$ and $N_{u}=20$, with the number of training steps $K=50000$ in Algorithm 2. The initial learning rate is set to $\alpha_{0}=0.001$, following a step decay learning rate schedule with a multiplicative factor of $\tau=0.95$ after every $1000$ iterations. To evaluate the performance of the trained model, we generate a test dataset $D_{test}=\{(\bm{x}_{1},\bm{d}_{1}),(\bm{x}_{2},\bm{d}_{2}),\ldots,(\bm{x}_{N_{test}},\bm{d}_{N_{test}})\}$ with $N_{test}=100$, following the same data generation process used for the training dataset $D_{train}$. For each pair $(\bm{x}_{i},\bm{d}_{i})$ in the test dataset $D_{test}$, we can calculate the corresponding approximate posterior $\nu_{\lambda}(\bm{\bm{v}_{i}})$ directly from the trained conditional network $\mathcal{N}_{\lambda}(\bm{v})$.
(a) Truth
(b) CNF-iVI
(c) pCN
@@
Figure 13: The comparison of the approximate posterior obtained by CNF-iVI method and the pCN method. The estimated variance and covariance functions obtained by the pCN algorithm are drawn in blue solid line, and by CNF-iVI algorithm the red dashed line. (a): The background truth of $u$. (b): The mean of approximate posterior obtained by CNF-iVI method. (c): The mean of posterior obtained by pCN method. (d): The covariance function $c(x,y)$ on all the mesh point pairs $\{(x_{i},x_{i})\}^{N_{g}}_{i=1}$; (e): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+40})\}^{N_{g}-40}_{i=1}$; (f): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+80})\}^{N_{g}-80}_{i=1}$.
Next, we compare the approximate posteriors obtained from the conditional network with the results generated by the pCN algorithm. For each pCN algorithm, we perform $3\times 10^{6}$ sampling iterations. Due to the high computational cost of the pCN algorithm, we limit our comparison to the posteriors corresponding to three examples in test dataset. The ESS of the pCN samples of these three examples are $ESS_{p_{1}}=267$, $ESS_{p_{2}}=231$, and $ESS_{p_{3}}=258$, respectively.
The conditional network provides a rough estimate, leading to a significant discrepancy between the mean of the approximate posterior and the mean of the true posterior. Consequently, the covariance function generated by the conditional network also significantly differs from the true posterior covariance. In the next subsection, we will present the results obtained after retraining the model using the same test dataset.
(a) Truth
(b) CNF-iVI
(c) pCN

### v2

@@ 4.2 Steady-State Darcy Flow Equation
Figure 8: The comparison of approximation ability between functional planar flow and functional Sylvester flow. We use 5 layers for both flow models to approximate the posterior. (a)(d): The background truth of $u$. (b): The mean of approximate posterior obtained by functional Sylvester flow. (c): The variance function of approximate posterior obtained by functional Sylvester flow. (e): The mean function of approximate posterior obtained by functional planar flow. (f):The variance function of approximate posterior obtained by functional planar flow.
4.3 Electrical Impedance Tomography
In this subsection, we focus on the electrical impedance tomography (EIT) problem. EIT is a technique to see inside an object by measuring electricity on its surface. It works by attaching electrodes, sending a small electrical current through them, and then reading the voltages that appear. We utilize the complete electrode model (CEM), proposed in somersalo1992existence. The strong form of the PDE governing this model is given as follows:
$\begin{cases}-\nabla\cdot(\kappa\nabla v(x))=0,&x\in\Omega,\\ \int_{e_{l}}\kappa\frac{\partial v}{\partial n}dS=I_{l},&l=1,\ldots,L,\\ \kappa\frac{\partial v}{\partial n}(x)=0,&x\in\partial\Omega\setminus\bigcup_{l=1}^{L}e_{l},\\ v(x)+z_{l}\kappa\frac{\partial v}{\partial n}(x)=V_{l},&x\in e_{l},l=1,\ldots,L,\end{cases}$
where $\Omega\subset\mathbb{R}^{2}$ is the domain and $\{e_{l}\}_{l=1}^{L}\subset\partial\Omega$ are electrodes on the boundary upon which currents $I_{\textbf{all}}=\{I_{l}\}_{l=1}^{L}$ are injected and voltages $\{V_{l}\}_{l=1}^{L}$ are read. The constants $\{z_{l}\}_{l=1}^{L}$ represent the contact impedances of the electrodes. The function $\kappa\in\mathcal{H}_{\kappa}$ represents the conductivity of the body and $v\in\mathcal{H}_{v}$ represents the potential within the body.
The inverse problem we consider is the recovery of $\kappa\in\mathcal{H}_{\kappa}$ from a sequence of boundary voltage measurements on the electrodes. The mathematical formulation of the inverse problem is given as follows.
Following somersalo1992existence, to improve the recovery of the model parameter $\kappa$, a standard strategy is to apply multiple distinct currents injection $I_{\mathrm{all}}^{(1)},I_{\mathrm{all}}^{(2)},\dots,I_{\mathrm{all}}^{(M)}$ and record the corresponding electrode voltages, thereby collecting richer measurement information. Under a single stimulation pattern $I_{\mathrm{all}}^{(i)}=\{I^{(i)}_{l}\}_{l=1}^{L}$, the solution of the PDE consists of the electric potential field $v^{(i)}$ inside the conductor and the set of voltage data $\{V_{l}^{(i)}\}_{l=1}^{L}$ on the $L$ electrodes, respectively. Thus, the PDE solution operator $\mathcal{G}^{(i)}$: $\mathcal{H}_{\kappa}\rightarrow\mathcal{H}_{v}\times\mathbb{R}^{L}$ can be written as
$\mathcal{G}^{(i)}(\kappa)=\left(v^{(i)},V^{(i)}_{1},V^{(i)}_{2},\ldots,V^{(i)}_{L}\right)$.
In practice, direct measurement of the internal voltage $v^{(i)}$ is challenging, and only the electrode voltages $\{V^{(i)}_{l}\}_{l=1}^{L}$ are accessible. Hence the measurement operator $\mathcal{S}^{(i)}$: $\mathcal{H}_{v}\times\mathbb{R}^{L}\rightarrow\mathbb{R}^{L}$ is defined as
$\mathcal{S}^{(i)}\left(v^{(i)},V^{(i)}_{1},V^{(i)}_{2},\ldots,V^{i}_{L}\right)=\left(V^{(i)}_{1},V^{(i)}_{2},\ldots,V^{(i)}_{L}\right)$.
We perform $M$ independent current stimulations, and the complete set of measured electrode voltages is therefore
$V=\left(V^{(1)},V^{(2)},\dots,V^{(M)}\right),$
where $V^{(m)}=\{V_{l}^{(m)}\}_{l=1}^{L}$ is the boundary voltage data of the $m$-th current stimulation, m=1,,Mm=1,...,M. Since $M$ current stimulations are performed, the inverse problem can be written as
$\bm{d}=\mathcal{S}\mathcal{G}(\kappa)+\bm{\epsilon},$
where $\bm{\epsilon}$ is the noise, $\mathcal{G}$ is a collection of PDE solution operators mapping from $\kappa$ to their corresponding solutions:
$\mathcal{G}(\kappa)=\left(\mathcal{G}^{(1)}(\kappa),\mathcal{G}^{(2)}(\kappa),\ldots,\mathcal{G}^{(M)}(\kappa)\right),$
and $\mathcal{S}$ is the measurement operator:
$\mathcal{S}\mathcal{G}(\kappa)=\left(\mathcal{S}^{(1)}\mathcal{G}^{(1)}(\kappa),\mathcal{S}^{(2)}\mathcal{G}^{(2)}(\kappa),\ldots,\mathcal{S}^{(M)}\mathcal{G}^{(M)}(\kappa)\right).$
(a) mesh
(b) measurement data
(c) measurement data
diag
Figure 9: (a) Mesh and electrodes. (b)(c) Measurement data $\bm{d}$ with noise $\bm{\epsilon}$.
For the numerical experiments conducted in this study, we specify the computational domain as $\Omega=\{x\in\mathbb{R}^{2}\mid\|x\|<1\}$, where $\|\cdot\|$ denotes the Euclidean norm in $\mathbb{R}^{2}$. Sixteen electrodes are equally spaced along the boundary $\partial\Omega$, achieving a 50% boundary coverage rate (see Subfigure (a) of Figure 9). All contact impedances are taken to be $z_{l}=0.025$. Adjacent electrodes are stimulated with a current of 1, so that the matrix of stimulation patterns $I_{\textbf{total}}\in\mathbb{R}^{16\times 15}$ ($I_{\textbf{total}}(l,i)=I_{l}^{(i)}$) is given by
$I=1\times\begin{pmatrix}+1&0&\cdots&0\\ -1&+1&\cdots&0\\ 0&-1&\ddots&0\\ \vdots&\vdots&\ddots&+1\\ 0&0&0&-1\end{pmatrix}$.
We assume the measurement data (Subfigures (b) and (c) if Figure 9) are generated from the underlying ground truth, depicted in Subfigures (a) and (d) of Figure 10. Since the parameter of interest in our model is a piecewise-constant function, we select the widely adopted level–set method burger2001level, iglesias2016bayesian, burger2005survey to solve the inverse problem. Simply put, we reparameterize the conductivity $\kappa$ via a level-‑set function $\mathcal{F}\mathrel{\mathop{\ordinarycolon}}\mathbb{R}\to\mathbb{R}$ as
$\kappa(x)=\mathcal{F}\bigl(\hat{\kappa}(x)\bigr),\qquad\mathcal{F}(x)=\begin{cases}1,&x\leq 0,\\[4.0pt] 3,&x>0.\end{cases}$
Consequently, the forward model becomes
$\bm{d}=\mathcal{S}\mathcal{G}\!\bigl(\mathcal{F}(\hat{\kappa})\bigr)+\bm{\epsilon},$
@@ 4.3 Electrical Impedance Tomography
A key limitation of the level-set method for reconstructing piecewise constant functions lies in its requirement that the exact function values are known a priori. For instance, in our EIT problem, the level-set function is constructed by predefining the constant values as 1 and 3. This simplification (i.e., assuming the function values are known prior to solving the inverse problem) is also extensively adopted in related studies dunlop2017hierarchical, iglesias2016bayesian. When assuming the constant values are unknown, more sophisticated techniques must be incorporated (see, e.g., akyildiz2025efficient), which fall beyond the main scope of this work. We thus merely acknowledge their relevance herein without providing further elaboration. Consistent with this simplification, our analysis presupposes that the constant values of the piecewise constant function are known a priori to solving the inverse problem.
For clarity, we list the specific choices for some parameters introduced in this problem as follows:
  • 1.
[…]
  • 2.
    The covariance operator $\mathcal{C}_{0}$ associated with the prior measure $\mu_{0}$ is defined as $\mathcal{C}_{0}=(\text{I}-\alpha\Delta)^{-2}$, where $\alpha=0.1$ denotes a pre-specified constant. Here, the Laplace operator $\Delta$ is defined on the computational domain $\Omega$ subject to homogeneous Neumann boundary conditions. Furthermore, the prior measure $\mu_{0}$ is taken to have a zero mean.
  • 3.
    To avoid inverse crime, the synthetic data is generated on a fine mesh consisting of $20054$ grid points, while a mesh of distinct resolution ($n=5124$ grid points) is adopted in the inversion stage.
@@
(b) mean (Sylvester)
(c) variance (Sylvester)
@@
(e) mean (Sylvester)
(f) variance (Sylvester)
Figure 10: (a)(d) Background truth. (b)(e) Mean of the approximate posterior. (c)(f) Pointwise variance of the approximate posterior.
Owing to the nonlinear nature of the inverse problem, we employ a five-layer functional Sylvester flow model to approximate the posterior distribution. Following Algorithm 1, the flow model is trained for $K=2000$ iterations, with each iteration drawing $N=10$ samples. A step-decay learning rate schedule is adopted: the initial learning rate is set to $\alpha_{0}=0.01$, and it is multiplied by a decay factor $\tau=0.8$ every 200 iterations. The application of the pCN algorithm to the EIT problem resulted in prohibitive computational cost. Consequently, solving the EIT problem with the pCN algorithm was omitted in our experiment.
In Subfigures (b), (c), (e), and (f) of Figure 10, we display the posterior measures generated by the functional Sylvester flow. The mean functions of the approximate posteriors are observed to closely resemble the true functions, as evidenced in Subfigures (a)(b) and (d)(e). Furthermore, Subfigures (c) and (f) depict the covariance function of the estimated posteriors. Collectively, these numerical experiments demonstrate the efficacy of our proposed algorithm for the EIT problem.
@@ 5 Numerical Examples of Conditional Functional Normalizing Flow
In this section, we present the results of the conditional functional normalizing flow, as described in Section 3. For a given measurement pair $(\bm{x}^{*},\bm{d}^{*})$, the trained network $\mathcal{N}_{\lambda}(\bm{v})$ can efficiently provides a rough estimate of the corresponding posterior distribution $\mu_{(\bm{x}^{*},\bm{d}^{*})}$. In addition, for a specific measurement pair, the results can be further refined through a retraining process.
@@ 5.1 Conditional Functional Normalizing Flow
In this subsection, we detail the training process of the conditional network and the extraction of an initial posterior approximation directly from it. Considering the steady-state Darcy flow equation (4.7), the inverse problem aims to infer the posterior distribution of the parameter $u$, given measurement data $\bm{d}=(w(x^{1}),\ldots,w(x^{N_{d}}))$ at points $\bm{x}=(x^{1},\ldots,x^{N_{d}})$. To approximate the posterior, we employ a 5-layer functional Sylvester flow similar with Subsection 4.2.
To effectively train the conditional network, we generate the training dataset based on the following process:
@@
    Draw a large number of samples from the measure $\mu_{0}$, denoted as $\{u_{1},u_{2},\ldots,u_{N_{train}}\}$, where $\mu_{0}$ is the prior measure we defined in Subsection 4.2. Examples of the samples drawn from $\mu_{0}$ are shown in Figure 11.
@@
(a) sample 1
(b) sample 2
(c) sample 3
diag
Figure 11: Samples from $\mu_{0}$.
Here, we briefly describe the symbols used in this context. The vector $\bm{x}=(x^{1},x^{2},\ldots,x^{\bm{N}_{d}})$ represents the measurement points, where each $x^{i}$ corresponds to a distinct measurement point. The notation $\bm{x}_{j}$ refers to different measurement points associated with different measurement pairs $(\bm{x}_{j},\bm{d}_{j})$. For convenience, all the measurement points $\bm{x}_{1},\ldots,\bm{x}_{N_{\text{train}}}$ are selected uniformly across the domain. Their coordinates are given by $\{(i/21,j/21)\}_{i,j=1}^{20}$.
(a) background truth
(b) mean (CNF-iVI)
(c) mean (pCN)
@@
Figure 12: The comparison of the measures obtained by CNF-iVI method and the pCN method. The estimated variance and covariance functions obtained by the pCN algorithm are drawn in blue solid line, and by CNF-iVI algorithm the red dashed line. (a): The background truth of $u$. (b): The mean of approximate posterior obtained by CNF-iVI method. (c): The mean of posterior obtained by pCN method. (d): The covariance function $c(x,y)$ on all the mesh point pairs $\{(x_{i},x_{i})\}^{N_{g}}_{i=1}$; (e): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+40})\}^{N_{g}-40}_{i=1}$; (f): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+80})\}^{N_{g}-80}_{i=1}$.
The model is trained on the dataset $D_{\text{train}}$ with Algorithm 2. We set $M=10$ and $N_{u}=20$, with the number of training steps $K=50000$ in Algorithm 2. The initial learning rate is set to $\alpha_{0}=0.001$, following a step decay learning rate schedule with a multiplicative factor of $\tau=0.95$ after every $1000$ iterations. To evaluate the performance of the trained model, we generate a test dataset $D_{test}=\{(\bm{x}_{1},\bm{d}_{1}),(\bm{x}_{2},\bm{d}_{2}),\ldots,(\bm{x}_{N_{test}},\bm{d}_{N_{test}})\}$ with $N_{test}=100$ following the same data generation process with training dataset $D_{train}$. For each pair $(\bm{x}_{i},\bm{d}_{i})$ in the test dataset $D_{test}$, the corresponding approximate posterior $\nu_{\lambda}(\bm{\bm{v}_{i}})$ can be computed directly using the trained conditional network $\mathcal{N}_{\lambda}(\bm{v})$.
(a) background truth
(b) mean (CNF-iVI)
(c) mean (pCN)
@@
Figure 13: The comparison of the measures obtained by CNF-iVI method and the pCN method. The estimated variance and covariance functions obtained by the pCN algorithm are drawn in blue solid line, and by CNF-iVI algorithm the red dashed line. (a): The background truth of $u$. (b): The mean of approximate posterior obtained by CNF-iVI method. (c): The mean of posterior obtained by pCN method. (d): The covariance function $c(x,y)$ on all the mesh point pairs $\{(x_{i},x_{i})\}^{N_{g}}_{i=1}$; (e): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+40})\}^{N_{g}-40}_{i=1}$; (f): The covariance function $c(x,y)$ on the mesh points $\{(x_{i},x_{i+80})\}^{N_{g}-80}_{i=1}$.
We now compare the approximate posterior obtained from the conditional network with that produced by the pCN algorithm. For each pCN run, we perform $3\times 10^{6}$ sampling iterations. Owing to the high computational cost of the pCN algorithm, we restrict our comparison to the posteriors corresponding to two examples in the test dataset.
The conditional network yields only a rough estimate, resulting in a substantial discrepancy between the mean of the approximate posterior and that obtained from the pCN algorithm. Consequently, the covariance function produced by the conditional network also differs significantly from that generated by the pCN algorithm. Figures 12 and 13 present qualitative results of the approximate posterior inferred using the trained conditional network $\mathcal{N}_{\lambda}(\bm{v})$ for two different measurement pairs. To further demonstrate the effectiveness of the CNF-iVI method, we compute the re-simulation error:
$\begin{split}\textbf{re-simulation error:}&=E_{(\bm{x},\bm{d})}\left(E_{\nu_{\lambda}(\bm{v}^{(\bm{x},\bm{d})})}\lVert\mathcal{S}\mathcal{G}(u)-\mathcal{S}\mathcal{G}(u_{truth}^{(\bm{x},\bm{d})})\rVert\right)\\ &\approx\frac{1}{N_{test}N_{samples}}\sum\limits_{i=1}^{N_{test}}\sum\limits_{j=1}^{N_{samples}}\lVert\mathcal{S}\mathcal{G}(u_{ij})-\mathcal{S}\mathcal{G}(u_{truth}^{(\bm{x}_{i},\bm{d}_{i})})\rVert\\ &=0.10834.\end{split}$ (5.1)
where $u_{truth}^{(\bm{x},\bm{d})}$ represents the ground truth and $\bm{v}^{(\bm{x},\bm{d})}$ denotes the vector corresponding to the measurement information $(\bm{x},\bm{d})$, and $N_{samples}=500$ for this evaluation.
The visual evidence presented in Figures 12 and 13, together with the quantitative results provided by the re-simulation error in (5.1), demonstrates the effectiveness of the proposed method in roughly capturing the properties of the posterior conditioned on the measurement information $(\bm{x},\bm{d})$. This further highlights the efficiency of the conditional functional normalizing flow in providing a rough inference of the posterior.
(a) background truth
(b) mean (retrained model)
(c) mean (pCN)